How do i solve x/10=4/5
Found 2 solutions by Alan3354, jfree12:
Answer by Alan3354(69443)

How do i solve x/10=4/5
-----------
Eliminate fractions by multiplying by 10

Answer by jfree12(1)

x/10=4/5 when solving proportions, you cross multiply. therefore x/10=4/5 you would come up with 5x=40, therefore x=8. plug 8 in for x and you get, 8/10=4/5. which is correct.